# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v1

@@ 1 Introduction
The mixture of experts (MoE) [15, 17] architecture has recently become a powerful approach in conditional computation, driving many advances in machine learning. Unlike dense models, MoEs dynamically activates only a subset of network, known as “expert”, for each input. This results in an efficient form of conditional computation. A notable modern example is the sparsely gated MoE [36], which significantly increases the model capacity without a corresponding increase in computational costs [19, 10, 11]. This has made MoEs crucial for scaling up large language [16, 30, 42, 8], vision [32, 20, 33], and multimodal models [24, 12], resulting in outstanding performance in different areas.
The mixture of experts (MoE) model includes $N$ expert networks and a gating network. Each expert ${\bm{h}}_{i}(\cdot;{\bm{\eta}}_{i})$ transforms the input ${\bm{x}}\in\mathbb{R}^{d}$ into an output in $\mathbb{R}^{d_{\rm o}}$. The gating network computes the gating probabilities ${\bm{g}}({\bm{x}};\bm{\Theta}_{g})=(g_{1}({\bm{x}};\bm{\Theta}_{g}),\ldots,g_{N}({\bm{x}};\bm{\Theta}_{g}))$, with $\bm{\Theta}_{g}$ as learnable parameters. The model output ${\bm{y}}$ is the weighted sum of expert outputs: ${\bm{y}}=\sum_{i=1}^{N}g_{i}({\bm{x}};\bm{\Theta}_{g})\cdot{\bm{h}}_{i}({\bm{x}};{\bm{\eta}}_{i})$. In the original MoE framework [15, 17], the gating network ${\bm{g}}$ computes scores for each expert using a parameterized network ${\bm{s}}({\bm{x}};\bm{\Theta}_{g})=\left(s_{1}({\bm{x}};{\bm{\theta}}_{1}),\ldots,s_{N}({\bm{x}};{\bm{\theta}}_{N})\right)\in\mathbb{R}^{N}$ and uses a softmax function to normalize these scores into a gating probability vector. Consequently, the scalar $g_{i}({\bm{x}};\bm{\Theta}_{g})$ can be expressed as
@@
The scoring network typically utilizes a linear mapping, which has been consistently adopted in modern sparse MoEs [36, 19, 10, 16, 6] due to its simplicity and scalability.
@@
Organization. The paper proceeds as follows. In Section 2, we provide background on MoE layer, attention mechanism and their connection. Next, we investigate the convergence behavior of parameter and expert estimation under the MoE model with two variants of the quadratic gate, and present two important practical implications from that analysis in Section 3. Subsequently, we conduct extensive experiments to empirically justify the theoretical results and favorable performance of the active-attention mechanism. Finally, we conclude the paper in Section 6. Full proofs and the remaining materials are deferred to the Appendices.
@@ 3.1 Quadratic Polynomial Gate
where $({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},c^{*}_{i},{\bm{\eta}}^{*}_{i})_{i=1}^{N^{*}}$ are unknown ground-truth parameters in $\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}$, and $G_{*}:=\sum_{i=1}^{k_{*}}\exp(c^{*}_{i})\delta_{({\bm{A}}_{i}^{*},{\bm{b}}^{*}_{i},{\bm{\eta}}^{*}_{i})}$ denotes the associated mixing measure, a weighted sum of Dirac measures $\delta$. Meanwhile, the function $h({\bm{x}};{\bm{\eta}})$ is referred to as the expert function, which we assumed to be of parametric form.
@@
We say that an expert function $x\mapsto h({\bm{x}},{\bm{\eta}})$ is strongly identifiable if it is twice differentiable w.r.t its parameter ${\bm{\eta}}$, and if for any $k\geq 1$ and pair-wise different parameters ${\bm{\eta}}_{1},\ldots,{\bm{\eta}}_{k}$, the following set
@@
Poly-strongly identifiable experts. To capture the convergence behavior of strongly identifiable experts, let us construct a loss function among parameters based on a notion of Voronoi cells [22, 26, 25]. Given an arbitrary mixing measure $G$ with $k^{\prime}\leq k$ components, we distribute its components to the following Voronoi cells, which are generated by the components of $G_{*}$:
$\mathcal{V}_{j}\equiv\mathcal{V}_{j}(G)$ :=$:=\{i\in[k^{\prime}]:\|\omega_{i}-\omega^{*}_{j}\|\leq\|\omega_{i}-\omega^{*}_{\ell}\|,\forall\ell\neq j\},$ (16)
@@
Suppose that the expert function $h(\cdot,{\bm{\eta}})$ is strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{k}(\Theta)$:
@@
There are two main implications from the result of Theorem 3.3. First, it follows from the formulation of the loss function $\mathcal{L}_{1}$ that exact-specified parameters ${\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j}$, i.e. $j\in[N^{*}]:|\mathcal{V}_{j}(\widehat{G}_{n})|=1$, share the same estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$. Note that as the expert $h(\cdot,{\bm{\eta}})$ is a Lipschitz function, then by denoting $\widehat{G}_{n}:=\sum_{i=1}^{\widehat{k}_{n}}\exp(\widehat{c}^{n}_{i})\delta_{(\widehat{{\bm{A}}}^{n}_{i},\widehat{{\bm{b}}}^{n}_{i},\widehat{{\bm{\eta}}}^{n}_{i})}$, we get
@@ 4 Practical Implications
where $\bar{\sigma}(.)$ is a non-linear function. We name the new attention mechanism (20) as active-attention. Our experiments with the active-attention in Figure 2 and Table 2 for both classification and time series forecasting tasks with a wide range of non-linear function $\bar{\sigma}(.)$ demonstrate the favorable performance of active-attention over the standard attention mechanism.
@@ 6 Discussion
In this supplementary material, we first discuss related works to the quadratic gating MoE model in Section A. Then, in Section C, we provide a convergence analysis for parameter and expert estimation under the quadratic gating mixture of linear experts. Full proofs for the theoretical results of Section 3 and Appendix C are presented in Appendix D. Next, we study the identifiability of the quadratic gating MoE in Appendix E. Lastly, we specify the experimental details for Figure 1 in Appendix F.
@@ Appendix B Quadratic Monomial Gate
$\mathcal{L}_{3}(G,G_{*}):=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|^{2}+\|\Delta{\bm{\eta}}_{ij}\|^{2}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|+\|\Delta{\bm{\eta}}_{ij}\|\Big{]}$
@@
Assume that the expert function $h({\bm{x}},{\bm{\eta}})$ is strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{k}(\Theta)$:
@@ Appendix D Proof of Theoretical Results
For the proof of the theorem, we first introduce some notation. Firstly, we denote by $\mathcal{F}_{k}(\Theta)$ the set of conditional densities of all mixing measures in $\mathcal{G}_{k}(\Theta)$, that is, $\mathcal{F}_{k}(\Theta):=\{f_{G}({\bm{x}}):G\in\mathcal{G}_{N}(\Theta)\}$. Additionally, for each $\delta>0$, the $L_{2}(\mu)$ ball centered around the regression function $f_{G_{*}}({\bm{x}})$ and intersected with the set $\mathcal{F}_{k}(\Theta)$ is defined as
$\mathcal{F}_{k}(\Theta,\delta):=\left\{f\in\mathcal{F}_{k}(\Theta):\|f-f_{G_{*}}\|_{L_{2}(\mu)}\leq\delta\right\}.$
In order to measure the size of the above set, Geer et. al. [38] suggest using the following quantity:
$\mathcal{J}_{B}(\delta,\mathcal{F}_{k}(\Theta,\delta)):=\int_{\delta^{2}/2^{13}}^{\delta}H_{B}^{1/2}(t,\mathcal{F}_{k}(\Theta,t),\|\cdot\|_{L_{2}(\mu)})~{}\mathrm{d}t\vee\delta,$ (28)
where $H_{B}(t,\mathcal{F}_{k}(\Theta,t),\|\cdot\|_{L_{2}(\mu))}$ stands for the bracketing entropy [38] of $\mathcal{F}_{k}(\Theta,u)$ under the $L^{2}$-norm, and $t\vee\delta:=\max\{t,\delta\}$. By using the similar proof argument of Theorem 7.4 and Theorem 9.2 in [38] with notations being adapted to this work, we obtain the following lemma:
@@
Take $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{F}_{k}(\Theta,\delta))$ that satisfies $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Then, for some universal constant $c$ and for some sequence $(\delta_{n})$ such that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$, we achieve that
$\mathbb{P}\Big{(}\|f_{\widehat{G}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}>\delta\Big{)}\leq c\exp\left(-\frac{n\delta^{2}}{c^{2}}\right),$
@@
$H_{B}(\varepsilon,\mathcal{F}_{k}(\Theta),\|.\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon),$ (29)
for any $20<\varepsilon\leq 1/$ 2. Indeed, for any function $f_{G}\in\mathcal{F}_{k}(\Theta)$, since the expert functions are bounded, we obtain that $f_{G}({\bm{x}})\leq M$ for almost everywhere ${\bm{x}}$, where $M$ is bounded constant of the expert functions. Let $\tau\leq\varepsilon$ and $\{\pi_{1},\ldots,\pi_{N}\}$ be the $\tau$-cover under the $L^{2}$ norm of the set $\mathcal{F}_{k}(\Theta)$ where $N:={N}(\tau,\mathcal{F}_{k}(\Theta),\|\cdot\|_{L^{2}(\mu)})$ is the ${\bm{\eta}}$-covering number of the metric space $(\mathcal{F}_{k}(\Theta),\|\cdot\|_{L^{2}(\mu)})$. Then, we construct the brackets of the form $[L_{i}({\bm{x}}),U_{i}({\bm{x}})]$ for all $i\in[N]$ as follows:
$L_{i}({\bm{x}})$ $:=\max\{\pi_{i}({\bm{x}})-\tau,0\},$
$U_{i}({\bm{x}})$ $:=\max\{\pi_{i}({\bm{x}})+\tau,M\}.$
From the above construction, we can validate that $\mathcal{F}_{k}(\Theta)\subset\cup_{i=1}^{N}[L_{i}({\bm{x}}),U_{i}({\bm{x}})]$ and $U_{i}({\bm{x}})-L_{i}({\bm{x}})\leq 2\min\{2\tau,M\}$. Therefore, it follows that
$\|U_{i}-L_{i}\|_{L_{2}(\mu)}^{2}=\int(U_{i}-L_{i})^{2}\mathrm{d}\mu({\bm{x}})\leq\int 16\tau^{2}\mathrm{d}\mu({\bm{x}})=16\tau^{2},$
@@
$H_{B}(4\tau,\mathcal{F}_{k}(\Theta),\|\cdot\|_{L_{2}(\mu)})\leq\log N=\log{N}(\tau,\mathcal{F}_{k}(\Theta),\|\cdot\|_{L^{2}(\mu)}).$ (30)
@@
Therefore, we need to provide an upper bound for the covering number $N$. In particular, we denote $\Delta:=\{(A,b,c)\in\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}:(A,b,c,{\bm{\eta}})\in\Theta\}$ and $\Omega:=\{{\bm{\eta}}\in\mathbb{R}^{q}:({\bm{A}},{\bm{b}},c,{\bm{\eta}})\in\Theta\}$. Since $\Theta$ is a compact set, $\Delta$ and $\Omega$ are also compact. Therefore, we can find $\tau$-covers $\Delta_{\tau}$ and ${\Omega}_{\tau}$ for $\Delta$ and $\Omega$, respectively. We can check that
$|\Delta_{\tau}|\leq\mathcal{O}_{P}(\tau^{-(d^{2}+d+1)k}),\quad|\Omega_{\tau}|\lesssim\mathcal{O}_{P}(\tau^{-qk}).$
@@
For each mixing measure $G=\sum_{i=1}^{k}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})}\in\mathcal{G}_{k}(\Theta)$, we consider other two mixing measures:
$\check{G}:=\sum_{i=1}^{k}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},\overline{{\bm{\eta}}}_{i})},\qquad\overline{G}:=\sum_{i=1}^{k}\exp(\overline{c}_{i})\delta_{(\overline{A}_{i},\overline{b}_{i},\overline{{\bm{\eta}}}_{i})}.$
@@
Here, $\overline{{\bm{\eta}}}_{i}\in{\Omega}_{\tau}$ such that $\overline{{\bm{\eta}}}_{i}$ is the closest to ${\bm{\eta}}_{i}$ in that set, while $(\overline{A}_{i},\overline{b}_{i},\overline{c}_{i})\in\Delta_{\tau}$ is the closest to $({\bm{A}}_{i},{\bm{b}}_{i},c_{i})$ in that set. From the above formulations, we get that
$\|f_{G}-f_{\check{G}}\|_{L_{2}(\mu)}^{2}$ $=\int\Bigg{[}\sum_{i=1}^{k}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}A_{j}{\bm{x}}+(b_{j})^{\top}{\bm{x}}+c_{j})}\cdot[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{]}^{2}~{}\mathrm{d}\mu({\bm{x}})$
$\leq k\int\sum_{i=1}^{k}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{]}^{2}~{}\mathrm{d}\mu({\bm{x}})$
$\leq k\int\sum_{i=1}^{k}~{}[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})]^{2}~{}\mathrm{d}\mu({\bm{x}})$
$\leq k\int\sum_{i=1}^{k}~{}[L_{1}\cdot\|{\bm{\eta}}_{i}-\overline{{\bm{\eta}}}_{i}\|]^{2}~{}\mathrm{d}\mu({\bm{x}})$
$\leq(kL_{1}\tau)^{2},$

### v2

@@ 1 Introduction
The mixture of experts (MoE) [15, 17] architecture has recently become a powerful approach in conditional computation, driving many advances in machine learning. Unlike dense models, MoEs dynamically activates only a subset of network, known as “expert”, for each input. This results in an efficient form of conditional computation. A notable modern example is the sparsely gated MoE [38], which significantly increases the model capacity without a corresponding increase in computational costs [19, 10, 11, 31]. This has made MoEs crucial for scaling up large language [16, 32, 44, 8], vision [34, 20, 35], and multimodal models [24, 12], resulting in outstanding performance in different areas.
The mixture of experts (MoE) model includes $N$ expert networks and a gating network. Each expert ${\bm{h}}_{i}(\cdot;{\bm{\eta}}_{i})$ transforms the input ${\bm{x}}\in\mathbb{R}^{d}$ into an output in $\mathbb{R}^{d_{\rm o}}$. The gating network computes the gating probabilities ${\bm{g}}({\bm{x}};\bm{\Theta}_{g})=(g_{1}({\bm{x}};\bm{\Theta}_{g}),\ldots,g_{N}({\bm{x}};\bm{\Theta}_{g}))$, with $\bm{\Theta}_{g}$ as learnable parameters. The model output ${\bm{y}}$ is the weighted sum of expert outputs: ${\bm{y}}=\sum_{i=1}^{N}g_{i}({\bm{x}};\bm{\Theta}_{g})\cdot{\bm{h}}_{i}({\bm{x}};{\bm{\eta}}_{i})$.
In the original MoE framework [15, 17], the gating network ${\bm{g}}$ computes scores for each expert using a parameterized network ${\bm{s}}({\bm{x}};\bm{\Theta}_{g})=\left(s_{1}({\bm{x}};{\bm{\theta}}_{1}),\ldots,s_{N}({\bm{x}};{\bm{\theta}}_{N})\right)\in\mathbb{R}^{N}$ and uses a softmax function to normalize these scores into a gating probability vector. Consequently, the scalar $g_{i}({\bm{x}};\bm{\Theta}_{g})$ can be expressed as
@@
The scoring network typically utilizes a linear mapping, which has been consistently adopted in modern sparse MoEs [38, 19, 10, 16, 6, 25] due to its simplicity and scalability.
@@
Organization. The paper proceeds as follows. In Section 2, we provide background on MoE layer, attention mechanism and their connection. Next, we investigate the convergence behavior of parameter and expert estimation under the MoE model with two variants of the quadratic gate, and present two important practical implications from that analysis in Section 3. Then, we highlight some practical implications from our theory in Section 4. Subsequently, we conduct extensive experiments to empirically justify the theoretical results and favorable performance of the active-attention mechanism in Section 5. Finally, we conclude the paper in Section 6. Full proofs and the remaining materials are deferred to the Appendices.
@@ 3.1 Quadratic Polynomial Gate
where $({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},c^{*}_{i},{\bm{\eta}}^{*}_{i})_{i=1}^{N^{*}}$ are unknown ground-truth parameters in $\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}$, and $G_{*}:=\sum_{i=1}^{N_{*}}\exp(c^{*}_{i})\delta_{({\bm{A}}_{i}^{*},{\bm{b}}^{*}_{i},{\bm{\eta}}^{*}_{i})}$ denotes the associated mixing measure, a weighted sum of Dirac measures $\delta$. Meanwhile, the function $h({\bm{x}};{\bm{\eta}})$ is referred to as the expert function, which we assumed to be of parametric form.
@@
We say that an expert function $x\mapsto h({\bm{x}},{\bm{\eta}})$ is strongly identifiable if it is twice differentiable w.r.t its parameter ${\bm{\eta}}$, and if for any $N\geq 1$ and pair-wise different parameters ${\bm{\eta}}_{1},\ldots,{\bm{\eta}}_{N}$, the following set
@@
Poly-strongly identifiable experts. To capture the convergence behavior of strongly identifiable experts, let us construct a loss function among parameters based on a notion of Voronoi cells [22, 27, 26]. Given an arbitrary mixing measure $G$ with $N^{\prime}\leq N$ components, we distribute its components to the following Voronoi cells, which are generated by the components of $G_{*}$:
$\mathcal{V}_{j}\equiv\mathcal{V}_{j}(G)$ :=$:=\{i\in[N^{\prime}]:\|\omega_{i}-\omega^{*}_{j}\|\leq\|\omega_{i}-\omega^{*}_{\ell}\|,\forall\ell\neq j\},$ (16)
@@
Suppose that the expert function $h(\cdot,{\bm{\eta}})$ is strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{N}(\Theta)$:
@@
There are two main implications from the result of Theorem 3.3. First, it follows from the formulation of the loss function $\mathcal{L}_{1}$ that exact-specified parameters ${\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j}$, i.e. $j\in[N^{*}]:|\mathcal{V}_{j}(\widehat{G}_{n})|=1$, share the same estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$. Note that as the expert $h(\cdot,{\bm{\eta}})$ is a Lipschitz function, then by denoting $\widehat{G}_{n}:=\sum_{i=1}^{\widehat{N}_{n}}\exp(\widehat{c}^{n}_{i})\delta_{(\widehat{{\bm{A}}}^{n}_{i},\widehat{{\bm{b}}}^{n}_{i},\widehat{{\bm{\eta}}}^{n}_{i})}$, we get
@@ 4 Practical Implications
where $\bar{\sigma}(.)$ is a non-linear function. We name the new attention mechanism (20) as active-attention. Our experiments with the active-attention in Figure 2 and Table 2 for both classification and time series forecasting tasks with a wide range of non-linear function $\bar{\sigma}(\cdot)$ demonstrate the favorable performance of active-attention over the standard attention mechanism.
@@ 6 Discussion
In this supplementary material, we first discuss related works to the quadratic gating MoE model in Appendix A. Then, we establish the expert estimation rates under the quadratic monomial gating MoE model in Appendix B. Subsequently, in Appendix C, we provide a convergence analysis for parameter and expert estimation under the quadratic gating mixture of linear experts. Full proofs for the theoretical results of Section 3 and Appendix C are presented in Appendix D. Next, we study the identifiability of the quadratic gating MoE in Appendix E. Lastly, we specify the experimental details for Figure 1 in Appendix F.
@@ Appendix B Quadratic Monomial Gate
$\mathcal{L}_{3}(G,G_{*}):=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|^{2}+\|\Delta{\bm{\eta}}_{ij}\|^{2}\Big{]}$ $+\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|+\|\Delta{\bm{\eta}}_{ij}\|\Big{]}$
@@
Assume that the expert function $h({\bm{x}},{\bm{\eta}})$ is strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{N}(\Theta)$:
@@ Appendix D Proof of Theoretical Results
For the proof of the theorem, we first introduce some notation. Firstly, we denote by $\mathcal{F}_{N}(\Theta)$ the set of conditional densities of all mixing measures in $\mathcal{G}_{N}(\Theta)$, that is, $\mathcal{F}_{N}(\Theta):=\{f_{G}({\bm{x}}):G\in\mathcal{G}_{N}(\Theta)\}$. Additionally, for each $\delta>0$, the $L_{2}(\mu)$ ball centered around the regression function $f_{G_{*}}({\bm{x}})$ and intersected with the set $\mathcal{F}_{N}(\Theta)$ is defined as
$\mathcal{F}_{N}(\Theta,\delta):=\left\{f\in\mathcal{F}_{N}(\Theta):\|f-f_{G_{*}}\|_{L^{2}(\mu)}\leq\delta\right\}.$
In order to measure the size of the above set, Geer et al. [40] suggest using the following quantity:
$\mathcal{J}_{B}(\delta,\mathcal{F}_{N}(\Theta,\delta)):=\int_{\delta^{2}/2^{13}}^{\delta}H_{B}^{1/2}(t,\mathcal{F}_{N}(\Theta,t),\|\cdot\|_{L^{2}(\mu)})~{}\mathrm{d}t\vee\delta,$ (28)
where $H_{B}(t,\mathcal{F}_{N}(\Theta,t),\|\cdot\|_{L^{2}(\mu))}$ stands for the bracketing entropy [40] of $\mathcal{F}_{N}(\Theta,u)$ under the $L^{2}$-norm, and $t\vee\delta:=\max\{t,\delta\}$. By using the similar proof argument of Theorem 7.4 and Theorem 9.2 in [40] with notations being adapted to this work, we obtain the following lemma:
@@
Take $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{F}_{N}(\Theta,\delta))$ that satisfies $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Then, for some universal constant $c$ and for some sequence $(\delta_{n})$ such that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$, we achieve that
$\mathbb{P}\Big{(}\|f_{\widehat{G}_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}>\delta\Big{)}\leq c\exp\left(-\frac{n\delta^{2}}{c^{2}}\right),$
@@
$H_{B}(\varepsilon,\mathcal{F}_{N}(\Theta),\|.\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon),$ (29)
for any $20<\varepsilon\leq 1/$ 2. Indeed, for any function $f_{G}\in\mathcal{F}_{N}(\Theta)$, since the expert functions are bounded, we obtain that $f_{G}({\bm{x}})\leq M$ for almost everywhere ${\bm{x}}$, where $M>0$ is some bounded constant of the expert functions. Let $\tau\leq\varepsilon$ and $\{\xi_{1},\ldots,\xi_{k}\}$ be the $\tau$-cover under the $L^{\infty}$ norm of the set $\mathcal{F}_{N}(\Theta)$ where $k:=N(\tau,\mathcal{F}_{N}(\Theta),\|\cdot\|_{L^{\infty}})$ is the $\tau$-covering number of the metric space $(\mathcal{F}_{N}(\Theta),\|\cdot\|_{L^{\infty}})$. Then, we construct the brackets of the form $[L_{i}({\bm{x}}),U_{i}({\bm{x}})]$ for all $i\in[k]$ as follows:
$L_{i}({\bm{x}})$ $:=\max\{\xi_{i}({\bm{x}})-\tau,0\},$
$U_{i}({\bm{x}})$ $:=\max\{\xi_{i}({\bm{x}})+\tau,M\}.$
From the above construction, we can validate that $\mathcal{F}_{N}(\Theta)\subset\cup_{i=1}^{k}[L_{i}({\bm{x}}),U_{i}({\bm{x}})]$ and $U_{i}({\bm{x}})-L_{i}({\bm{x}})\leq\min\{2\tau,M\}$. Therefore, it follows that
$\|U_{i}-L_{i}\|_{L_{2}(\mu)}^{2}=\int(U_{i}-L_{i})^{2}\mathrm{d}\mu({\bm{x}})\leq\int 4\tau^{2}\mathrm{d}\mu({\bm{x}})=4\tau^{2},$
@@
$H_{B}(2\tau,\mathcal{F}_{N}(\Theta),\|\cdot\|_{L_{2}(\mu)})\leq\log k=\log N(\tau,\mathcal{F}_{N}(\Theta),\|\cdot\|_{L^{\infty}}).$ (30)
@@
Therefore, we need to provide an upper bound for the covering number $N(\tau,\mathcal{F}_{N}(\Theta),\|\cdot\|_{L^{\infty}})$. In particular, we denote $\Delta:=\{{\bm{A}},{\bm{b}},c)\in\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}:({\bm{A}},{\bm{b}},c,{\bm{\eta}})\in\Theta\}$ and $\Omega:=\{{\bm{\eta}}\in\mathbb{R}^{q}:({\bm{A}},{\bm{b}},c,{\bm{\eta}})\in\Theta\}$. Since $\Theta$ is a compact set, $\Delta$ and $\Omega$ are also compact. Therefore, we can find $\tau$-covers $\Delta_{\tau}$ and ${\Omega}_{\tau}$ for $\Delta$ and $\Omega$, respectively. We can check that
$|\Delta_{\tau}|\leq\mathcal{O}(\tau^{-(d^{2}+d+1)N}),\quad|\Omega_{\tau}|\lesssim\mathcal{O}(\tau^{-qN}).$
@@
For each mixing measure $G=\sum_{i=1}^{N}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})}\in\mathcal{G}_{N}(\Theta)$, we consider other two mixing measures:
$\check{G}:=\sum_{i=1}^{N}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},\overline{{\bm{\eta}}}_{i})},\qquad\overline{G}:=\sum_{i=1}^{N}\exp(\overline{c}_{i})\delta_{(\overline{{\bm{A}}}_{i},\overline{{\bm{b}}}_{i},\overline{{\bm{\eta}}}_{i})}.$
@@
Here, $\overline{{\bm{\eta}}}_{i}\in{\Omega}_{\tau}$ such that $\overline{{\bm{\eta}}}_{i}$ is the closest to ${\bm{\eta}}_{i}$ in that set, while $(\overline{{\bm{A}}}_{i},\overline{{\bm{b}}}_{i},\overline{c}_{i})\in\Delta_{\tau}$ is the closest to $({\bm{A}}_{i},{\bm{b}}_{i},c_{i})$ in that set. From the above formulations, we get that
$\|f_{G}-f_{\check{G}}\|_{L^{\infty}}$ $=\Bigg{\|}\sum_{i=1}^{N}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})]\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\Bigg{\|}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}\cdot[h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})]\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\|h({\bm{x}},{\bm{\eta}}_{i})-h({\bm{x}},\overline{{\bm{\eta}}}_{i})\|_{L^{\infty}}$
$\lesssim\sum_{i=1}^{N}\|{\bm{\eta}}_{i}-\overline{{\bm{\eta}}}_{i}\|\lesssim\tau.$